Define: R^3→R^3 by

T(x,y,z)=(x-y+z,x+y,y+z)

Let v1= (1,1,1), v2= (0,1,1), v3= (0,0,1). Find a matrix of T with respect to the basis {v1,v2,v3}. Futher check T is invertible or not.




1
Expert's answer
2021-02-23T10:16:40-0500

The map "T" with respect to the standard basis "e_1=(1,0,0)", "e_2=(0,1,0)", "e_3=(0,0,1)" can be presented in the form: "\left(\begin{matrix}x-y+z\\x+y\\y+z\end{matrix}\right)=\left(\begin{matrix}1&-1&1\\1 &1&0\\0&1&1\end{matrix}\right)\left(\begin{matrix}x\\y\\z\end{matrix}\right)". In order to find a matrix with respect to the basis "v_1","v_2" and "v_3" , we need to find such matrix "T'" that "x'v_1+y'v_2+z'v_3=T'(xv_1+yv_2+zv_3)". We have: "xv_1+yv_2+zv_3=(x,x+y,x+y+z)^{\top}" In case we act on the vector with the map "T" we will get: "(x-(x+y)+x+y+z,2x+y,2x+2y+z)=(x+z,2x+y,2x+2y+z)". It has the following representation in basis "v_1,v_2,v_3:" "(x+z,2x+y,2x+2y+z)=(x+z)v_1+(2x+y-(x+z))v_2+(2x+2y+z-(x+z)-(x+y-z))v_3=(x+z)v_1+(x+y-z)v_2+(y+z)v_3"

The respective matrix has the form: "T'=\left(\begin{matrix}1&0&1\\1&1&-1\\0&1&1\end{matrix}\right)"